OpenStudy (anonymous):

what is the missing term in this geometric sequence? ...,-32,_,-128,...

OpenStudy (solomonzelman):

\[-32 \times x^2=-128~~~~~~~x=common~~ratio,~~~~~x=?\]

OpenStudy (solomonzelman):

\[x^2=-128/-32=4~~~~x^2=4~~~So,~~x=+-2\] it's either \[-32 \times (-2)=64~~~~OR~~~~-32 \times (2)=-64\] So it's either 64 or -64.